OpenStudy (anonymous):

A baseball is thrown up into the air. The ball's height in feet is modeled by the equation h(t) = -16t2 + 32t + 10. (h = height, t = time in seconds) Find the maximum height of the ball.

OpenStudy (anonymous):

Differentiate h[t] h'(t)=-32t+32 at maxima h'(t)=0 32=32t t=1; so h max=-16+32+10=26.